ZERO ELECTRON KINETIC ENERGY PULSED FIELD IONIZATION (ZEKE) SPECTROSCOPY OF CdCH_{3} AND ZnCH_{3}

\underline{SERGEY I. PANOV}, DAVID E. POWERS, TIMOTHY A. BARCKHOLTZ, AND TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

The two-color (1 + 1′) REMPI and ZEKE spectra of the CdCH_{3} and ZnCH_{3} radicals are reported for the first time. Based upon the observed intensities in the ZEKE spectrum for CdCH_{3}, via transitions to different vibrational levels of the Ã^{2}E intermediate state, the band previously assigned as the upper spin-orbit component of the vibrationless level of the Ã^{2}E ← X̃^{2}A_{1} transition in neutral CdCH_{3} is reassigned. This band which has now been observed in both LIF and REMPI experiments, is reassigned as the 6_{0}^{1} band of the lower spin-orbit component, ^{2}E_{1/2}. The ZEKE spectra, taken via various intermediate rovibronic levels, are used to reveal rovibrational structure of the CdCH_{3} and ZnCH_{3} cations and to measure the adiabatic ionization energies for both radicals. These radicals belong to the C_{3v} point group and are therefore subject to a Jahn-Teller interaction in the Ã state, which allows a broader range of K′ states to be accessible in the LIF and REMPI experiments thereby enabling access to a broader range of K states for the ions in the ZEKE experiment. The resolved K structure in the level with one quanta of ν_{6} excitation in the ion enables estimates of both the rotational constant A and the coriolis coupling constant aζ_{t} to be made. Vibrational frequencies and assignments for the ions will be reported.